OpenStudy (hhopke):

8 equals the square root of 3x solve for x

OpenStudy (misty1212):

HI!!

OpenStudy (tkhunny):

\(8 = \sqrt{3x}\) or \(8 = \sqrt{3}x\)

OpenStudy (hhopke):

the first one

OpenStudy (tkhunny):

This is an example of an inadequate response. Please show YOUR work. Anything you have will be welcomed.

Elsa213 (elsa213):

+you listed no answer choices

OpenStudy (vheah):

\[8=\sqrt{3x}\] This is when you work backwords. To get rid of the square root you have to square it, and you will be squaring the 8 as well. \[8^{2} = (\sqrt{3x})^{2}\] and it leaves you with: \[64=3x\] Divide three on both sides: \[64/3 = 3x/3\] \[64/3 = x\]